you have 24 square pieces of wood, describe all the different ways you could make a rectangle bu placing them side to side.
Found 2 solutions by ikleyn, Theo:
Answer by ikleyn(52858)

.
you have 24 square pieces of wood, describe all the different ways you could make a rectangle bu placing them side to side.
-----------------------------------------------------

In how many ways you can split the number "24" into a product of two factors?

The same number.


24 = 1*24 = 2*12 = 3*8 = 4*6 = 6*4 = 8*3 = 12*2 = 24*1.


Answer. In 8 ways.


Answer by Theo(13342)

you have 24 square pieces of wood.

you could lay them out in the following fashion.

first number is the number of rows.
second number is the numbe of columns.
for example:
1,24 is 1 row by 24 columns.
here goes:
1,24
2,12
3,8
4,6
6,4
8,3
12,2
24,1

i believe that's all.
it winds up being the number of ways you can multiply two integers to get 24.

here's a couple of examles of what it would look like visually.

$$$